3 white boxes and 4 blue boxes weigh 36 pounds.4 white boxes and 3 blue boxes weigh 34 pounds . what is the weight of one white box and one blue box ?
a 4 b 6 c 10 d 12 e 24

Answer by checkley77(12844)

3W+4B=36 MULTIPLY BY 4
4W+3B=34 MULTIPLY BY -3 & ADD
12W+16B=144
-12W-9B=-102
------------------
7B=42
B=42/7
B=6 POUNDS FOR EACH BLUE BOX.
3W+4*6=36
3W+24=36
3W=36-24
3W=12
W=12/3
W=4 POUNDS FOR EACH WHITE BOX.
6+4=10 POUNDS FOR ONE OF EACH.
PROOF:
4*4+3*6=34
16+18=34
34=34
ANSWER: C).